Determine the amount to be paid in full settlement of each of two invoices, (a) and (b)



PE 6-4A 

Determine the amount to be paid in full settlement of each of two invoices, (a) and (b), assuming that credit for returns and allowances was received prior to payment and that all invoices were paid within the discount period.



Merchandise
Freight
Paid by Seller

Freight Terms
Returns and
Allowances
a.$90,000
$1,000

FOB shipping point, 1/10, n/30$15,000
b.110,000
1,575

FOB destination, 2/10, n/308,500


PE 6-4B 

Determine the amount to be paid in full settlement of each of two invoices, (a) and (b), assuming that credit for returns and allowances was received prior to payment and that all invoices were paid within the discount period.



Merchandise
Freight
Paid by Seller

Freight Terms
Returns and
Allowances
a.$36,000
$800

FOB shipping point, 1/10, n/30$4,000
b.44,900
375

FOB destination, 2/10, n/302,400
Answer #1

6 4- A

a b
Merchandise value 90,000 110,000
Less: Returns 15,000 8,500
Net purchases 75,000 101,500
Discount @1% / 2% 750 2030
Net purchases of discount 74,250 99,470
Add: Freight 1000
Amount to be paid in full settlement 75,250 99,470

6 4 B

a b
Merchandise value 36,000 44,900
Less: Returns 4,000 2,400
Net purchases 32,000 42,500
Discount @1% / 2% 320 850
Net purchases of discount 31,680 41,650
Add: Freight 0 375
Amount to be paid in full settlement 31,680 42,025
Answer #2
Concepts and reason

FOB (Free on Board or Freight on Board) shipping point: This means the seller sends the inventory “free on board”. The buyer bears the freight or transportation or shipment charges from the point of origin to the point of destination. In the case of FOB shipping point, the title of the ownership is transferred to the buyer from that point of origin.

FOB destination: When the seller bears the freight or transportation costs from the point of origin to the point of destination, it is referred as FOB destination. The seller does not transfer the title of the ownership to the buyer until the inventory reaches the delivery point.

Fundamentals

Purchases: Purchase is the expense amount paid to a company in exchange for acquiring the goods or services. Purchase of inventory is added to the beginning merchandise inventory to compute cost of goods available for sale.

Net purchases: The amount of gross purchases less returns, allowances, and discounts are known as net purchases. The following is the formula used for net purchases:

Netpurchases={GrosspurchasesPurchasereturnsandallowancesPurchasediscounts+Freightcharge}{\rm{Net purchases = }}\left\{ \begin{array}{l}\\{\rm{Gross purchases}}\\\\{\rm{ -- Purchase returns and allowances}}\\\\{\rm{-- Purchase discounts + Freight charge}}\\\end{array} \right\}

Merchandise inventory: Merchandise is the goods procured by a trader, a wholesaler, or a merchant from suppliers with an intention to resell those goods for a higher price. This inventory is recorded as current asset on the balance sheet because it is planned to be sold within a year.

Invoice: This is a document that lists the details of goods sent or services rendered and the statement of amount outstanding. This is simply also called as bill.

Purchase returns: Purchase returns are the amount of merchandise returned back to the seller due to reasons like quality failure.

Purchase allowance: The amount the selling company grants as an allowance to the customer for agreeing to accept the defective or damaged goods is known as purchase allowance.

Purchase discount: The seller offers a concession or reduction in the purchase price for early payment. This reduction is known as purchase discount. The following is the formula used to compute purchase discount:

Purchasediscount=((GrosspurchasesPurchasereturnsandallowances)×Purchasediscountpercentage){\rm{Purchase discount = }}\left( \begin{array}{l}\\\left( \begin{array}{l}\\{\rm{Gross purchases}}\\\\{\rm{-- Purchase returns and allowances}}\\\end{array} \right)\\\\{\rm{ \times Purchase discount percentage}}\\\end{array} \right)

a)

Calculate the total amount to be paid in full settlement of invoice:

Particulars
Merchandise value
Less: Returns and allowances
Net merchandise value
Less: Purchase discount ($75,000 < 1%
Mercha

Therefore, the total amount to be paid is $75,250.

b)

Calculate the total amount to be paid in full settlement of invoice:

Particulars
Merchandise value
Less: Returns and allowances
Net merchandise value
Less: Purchase discount ($101,500 x 2%)
Tota

Therefore, the total amount to be paid is $99,470.

Ans: Part a

The total amount to be paid is $75,250.

Part b

The total amount to be paid is $99,470.